# S-Duality for Non-Abelian Monopoles

### v1

@@ 1 Introduction
In [1], Goddard, Nuyts and Olive showed that magnetic charges of monopoles in a theory with unbroken gauge group $H$ are classified by the weight lattice of the GNO dual group $H^{\vee}$, and conjectured that $H$-monopoles transform as $H^{\vee}$-multiplets, with the true symmetry group given by $H\times H^{\vee}$. In the context of $\mathcal{N}=4$ super-Yang-Mills theory, this kinematical GNO duality is embedded into the Montonen-Olive S-duality, which relates a theory with gauge group $G$ to a dual theory with gauge group $G^{\vee}$ and inverted coupling [2, 3]. Under the symmetry breaking $G\to H$ and $G^{\vee}\to H^{\vee}$, the duality identifies monopoles in one theory with W-bosons in the dual theory. Including the $\theta$-angle, the S-duality group of $\mathcal{N}=4$ SYM is $SL(2,\mathbb{Z})$, under which BPS dyons are permuted and, in particular, are mapped to W-bosons [4, 5].
In $\mathcal{N}=4$ SYM, the low-energy dynamics of BPS monopoles is governed by a supersymmetric quantum mechanics on the monopole moduli space, in which each bosonic modulus is paired with two fermionic partners [6, 7]. Quantization of the center-of-mass sector produces the required $16$-fold degeneracy of the massive vector multiplet, while BPS ground states on the relative moduli space are realized as (normalizable) harmonic forms [8, 4]. In the maximally broken phase $G\to U(1)^{r}$, there are $r$ fundamental monopoles (one for each simple coroot), which, together with their BPS bound states, reproduce the W-boson spectrum in the dual theory [9, 10, 11, 12]. With the electric charge turned on, dyonic bound states built from these monopoles realize the $SL(2,\mathbb{Z})$ S-duality orbit of the W-bosons [4, 13].
@@
Our approach may be viewed as a smooth-monopole analogue of the Kapustin-Witten construction for geometric Langlands program [41], where ’t Hooft operators in the topologically twisted $\mathcal{N}=4$ theory are defined by imposing supersymmetric singular monopole behavior, and the resulting magnetic data are encoded in the geometry of Hecke modifications. In this framework, a ’t Hooft (Hecke) operator $T(w^{\vee})$ is labeled by a dominant weight $w^{\vee}$ of the Langlands dual group $G^{\vee}$, and the moduli space of singular monopoles inserted at a single point is identified with the space $\mathscr{Y}(w^{\vee})$ of Hecke modifications of type $w^{\vee}$ for a $G$-bundle. Its natural compactification $\overline{\mathscr{Y}}(w^{\vee})$ is a Schubert variety in the affine Grassmannian. The space of physical states $\mathcal{H}(w^{\vee})$ is given by the (intersection) cohomology of $\overline{\mathscr{Y}}(w^{\vee})$ and is isomorphic to $R(w^{\vee})$, the irreducible $G^{\vee}$-representation of highest weight $w^{\vee}$. In this paper we consider smooth monopoles arising from symmetry breaking, but the underlying geometric picture is similar. In particular, compactifying the Hecke modification space $\mathscr{Y}(w^{\vee})$ by adding all associated lower-weight Hecke modifications (which share the same topological type determined by $w^{\vee}$) mirrors the stratification of the monopole moduli space by allowing gauge-inequivalent magnetic charges at fixed topological charge.
@@
For example, when $SU(N+1)$ breaks to $U(N)$, there are NN $SU(2)$-embedded monopole solutions of unit topological charge, related by global $SU(N)$ Weyl transformations [33]. They are the classical counterparts of dual W-bosons in the fundamental of $SU(N)$. With (1.2) relaxed to (1.4), these solutions are naturally incorporated into a single connected moduli space labeled by the topological charge. The low-energy effective action for monopoles is obtained by a collective-coordinate expansion of $\mathcal{N}=4$ SYM. Starting from an arbitrary BPS trajectory, integrating out $A_{0}$ automatically enforces the background-gauge condition on the moduli derivatives. We will argue that, even with the background gauge imposed, the non-Abelian modes that do not commute with $G_{0}$ still have divergent norm. This divergence is also required for the stratified structure of the moduli space. For instance, in a moduli space consisting of an open stratum $\mathbb{C}^{2N}$ and a closed stratum $\mathbb{CP}^{2N-1}$, $\mathbb{CP}^{2N-1}$ must sit at the asymptotic boundary of $\mathbb{C}^{2N}$ and thus inherits a divergent metric. With the non-Abelian modes included, the moduli space is not necessarily hyper-Kähler. To be self-contained, we will derive the effective action via a collective-coordinate expansion without relying on hyper-Kähler structure [^2] The hyper-Kähler structure is needed only for the fermionic zero-mode counting, where it guarantees that the number of fermionic zero modes is twice the number of bosonic moduli. Here we take this relation as an input. In fact, once the non-normalizable modes are included, the index calculation [21, 22] yields a moduli-space dimension that agrees with the stratified formulation incorporating non-Abelian modes [38, 39]. Although this dimension is not necessarily divisible by $4$, the index still gives a $2:1$ ratio of fermionic to bosonic modes for $\mathcal{N}=4$ SYM [43, 44].. The resulting action takes the same form as in the maximally broken case, but the associated supersymmetry is $\mathcal{N}=2$ (enhanced to $\mathcal{N}=4$ when the moduli space is Kähler), rather than the $\mathcal{N}=8$ available for a hyper-Kähler target [42]. This reduction does not affect the physical conclusions: the $16$-fold degeneracy of the $\mathcal{N}=4$ vector multiplet is already saturated by the eight fermionic zero modes associated with the center-of-mass motion, while the ground states in the relative sector are still given by harmonic forms on the moduli space.
@@
Aside from the numerical match, the monopole ground states, which furnish the same $(H^{\vee})^{s}$-representation as the dual W-bosons, also carry definite weights of that representation (up to Weyl transformations). This allows us to construct magnetic operators generating the $(H^{\vee})^{s}$-action on the monopole Hilbert space. While the $H^{s}$-action is realized geometrically by isometric diffeomorphisms of the relative moduli space, the $(H^{\vee})^{s}$-action is realized algebraically on differential forms by wedge and contraction operations. In [20], it was noted that the harmonic forms on the coset space, which represent monopole ground states in that case, are $H^{s}$-invariant. Here we show that the ground states are $H^{s}$-invariant in general. Consequently, the $(H^{\vee})^{s}$-action may be implemented by operators commuting with both $H^{s}$ and the Hamiltonian. For monopole moduli spaces of topological charge $1$, we give an explicit construction of the resulting $H^{s}\times(H^{\vee})^{s}$-representation on the Hilbert space, thereby realizing the $H\times H^{\vee}$ GNO conjecture at the level of monopole quantum mechanics.
@@
The rest of the paper is organized as follows. Section 2 reviews the stratified structure of the monopole moduli space. Section 3 presents a general proof of the matching between the monopole (and dyon) spectrum and the dual W-boson spectrum when the unbroken gauge group is non-Abelian. In Section 4 we study the supersymmetric quantum mechanics on the monopole moduli space with the non-Abelian degrees of freedom included. In Section 5 we construct the magnetic gauge transformation generators and analyze their relation to the electric generators. We end with a discussion in Section 6.
@@ 2 Stratification of the monopole moduli space
Let $\bm{\alpha}_{a}^{\vee}=2\bm{\alpha}_{a}/(\bm{\alpha}_{a}\cdot\bm{\alpha}_{a})$ be the coroots. The unbroken gauge group is $H=H^{s}\times U(1)^{t}$, where $H^{s}$ is semisimple with simple roots $\bm{\alpha}_{t+1},\ldots,\bm{\alpha}_{r}$ (determined up to the Weyl group of $H^{s}$).
@@
$m_{a}$ are gauge invariant, whereas $q_{a}$ transform under the Weyl group of $H$. For each $1\leq i\leq n$, the intersection $\mathfrak{h}_{0}\cap C(k_{i})$ is a Weyl ($H$)-orbit. We can choose a unique representative by imposing the anti-dominant condition
@@
    $\mathcal{M}_{\mathrm{rel}}(1,\Phi_{0})=\mathcal{M}_{\mathrm{rel},1}(1,\Phi_{0})=C(k_{1})\cong\mathbb{CP}^{1},$ (2.25)
    with $\text{dim}\mathcal{M}_{\mathrm{rel}}(1,\Phi_{0})=2$.
@@
    The stratum $\mathcal{M}_{\mathrm{rel},2}(2,\Phi_{0})$ contains a parameter $a\in[0,\infty)$ describing the size of the non-Abelian cloud. The corresponding monopole solutions and the moduli space metric can be found in [45, 46, 47]. When $a\to\infty$, $G_{0}$ jumps from $k_{2}$ to $C(k_{1})$, consistent with the connectedness of $\mathcal{M}_{\mathrm{rel}}(2,\Phi_{0})$. Note that the boundary of $\mathcal{M}_{\mathrm{rel},2}(2,\Phi_{0})$ would be $7$-dimensional, whereas $\mathcal{M}_{\mathrm{rel},1}(2,\Phi_{0})$ is $6$-dimensional, so $\mathcal{M}_{\mathrm{rel}}(2,\Phi_{0})$ is not obtained by attaching a boundary to $\mathcal{M}_{\mathrm{rel},2}(2,\Phi_{0})$, but rather by compactifying $\mathcal{M}_{\mathrm{rel},2}(2,\Phi_{0})$ with $\mathcal{M}_{\mathrm{rel},1}(2,\Phi_{0})$ placed at infinity, analogous to compactifying $\mathbb{R}^{2}$ to $\mathbb{S}^{2}$ by adding a point at infinity.
@@ 3.2 $\mathcal{M}_{\mathrm{rel}}(1,\Phi_{0})=C(k_{1})\cup\mathcal{M}_{\mathrm{rel},2}(1,\Phi_{0})$
    where $S$ is the tautological rank-$2$ subbundle on $\mathrm{Gr}(2,N{+}1)$ and $\pi$ is the bundle projection. Gluing this open bundle to the Schubert divisor at infinity yields a compactification [48]
@@ 3.3 General proof
For a compact, connected semisimple group $\mathcal{G}$, let $\lambda$ be an anti-dominant coweight of $\mathcal{G}$ (equivalently, a weight of $\mathcal{G}^{\vee}$), and let $V_{\lambda}$ be the irreducible $\mathcal{G}^{\vee}$-representation with lowest weight $\lambda$. The set of weights in $V_{\lambda}$ decomposes into Weyl orbits
@@
      $H^{\ast}[\mathcal{M}_{\mathrm{rel},i}(1,\Phi_{0})]=H^{\ast}[C(k_{i})],$ (3.68)
@@
      $\text{dim}H^{\ast}[\mathcal{M}_{\mathrm{rel},i}(1,\Phi_{0})]=\chi[\mathcal{M}_{\mathrm{rel},i}(1,\Phi_{0})]=\chi[C(k_{i})].$ (3.69)
@@
    In the full moduli space $\mathcal{M}(m,\Phi_{0})$, when $m>1$, each fiber $\mathcal{M}(m,\Phi_{0},k_{i})$ contains $m$ identical massive monopoles (with some additional massless monopoles if $i\neq 1$) and thus has the structure [4, 25]
@@
    In fact, when $m>1$ with $\mathcal{M}_{\mathrm{rel}}(m,\Phi_{0})$ noncompact, bound states are normalizable harmonic forms, which are typically not captured by ordinary de Rham cohomology. In our situation, the metric scale along $C(k_{i})$ is infinite, so the usual normalizability criterion must be adapted to the fiber-base decomposition. Let $N_{\mathcal{M}}$ denote the number of bound states on $\mathcal{M}$. Instead of (3.75), a suitable counting is
@@
    and all topological charges become holomorphic except for mm. [^5] In the $SU(3)$ example 2 for $m=2$, take $\tilde{\Phi}_{0}=\mathop{\mathrm{diag}}\nolimits(2v,-v+\varepsilon,-v-\varepsilon)$ with $0<\varepsilon<3v$. Both $\mathcal{M}_{\mathrm{rel}}(\tilde{\Phi}_{0},k_{1})$ and $\mathcal{M}_{\mathrm{rel}}(2,\Phi_{0},k_{1})$ are the relative moduli space of two massive monopoles of identical charge $\mathop{\mathrm{diag}}\nolimits(1,-1,0)$, whereas $\mathcal{M}_{\mathrm{rel}}(\tilde{\Phi}_{0},k_{2})$ contains an additional monopole of charge $\mathop{\mathrm{diag}}\nolimits(0,1,-1)$. As $\varepsilon\to 0$, the $\mathop{\mathrm{diag}}\nolimits(0,1,-1)$ sector becomes massless and form the non-Abelian cloud in $\mathcal{M}_{\mathrm{rel}}(2,\Phi_{0},k_{2})$ [25]. Here no $\mathbf{k}_{i}$ is a root, so $N_{\mathcal{M}_{\mathrm{rel}}(2,\Phi_{0})}=0$. We may expect [^6] When $\tilde{\Phi}_{0}\rightarrow\Phi_{0}$, harmonic forms on $\mathcal{M}_{\mathrm{rel}}(\tilde{\Phi}_{0},k_{i})$ descend to $\mathcal{M}_{\mathrm{rel}}(m,\Phi_{0},k_{i})$. This limiting procedure has subtleties, see [25]. We return to these issues in Example 3.3.
@@
    In the maximally broken regime, matching the monopole spectrum to the W-bosons requires

### v2

@@ 1 Introduction
In [1], Goddard, Nuyts and Olive showed that magnetic charges of monopoles in a theory with unbroken gauge group $H$ are classified by the weight lattice of the GNO dual group $H^{\vee}$, and conjectured that $H$-monopoles transform as $H^{\vee}$-multiplets, with the true symmetry group given by $H\times H^{\vee}$. In the context of $\mathcal{N}=4$ super-Yang-Mills theory, this kinematical GNO duality is embedded into the Montonen-Olive S-duality, which relates a theory with gauge group $G$ to a dual theory with gauge group $G^{\vee}$ and inverted coupling [2, 3]. Under the symmetry breaking $G\to H$ and $G^{\vee}\to H^{\vee}$, the duality identifies monopoles in one theory with W-bosons in the dual theory. Including the $\theta$-angle, the S-duality group of $\mathcal{N}=4$ SYM is $SL(2,\mathbb{Z})$, under which monopoles and dyons are permuted and, in particular, are mapped to the dual W-bosons [4, 5].
In $\mathcal{N}=4$ SYM, the low-energy dynamics of BPS monopoles is governed by a supersymmetric quantum mechanics on the monopole moduli space, in which each bosonic modulus is paired with two fermionic partners [6, 7]. Quantization of the center-of-mass sector produces the required $16$-fold degeneracy of the massive vector multiplet, while BPS ground states on the relative moduli space are realized as (normalizable) harmonic forms [8, 4]. In the maximally broken phase $G\to U(1)^{r}$, there are $r$ fundamental monopoles (one for each simple coroot), which, together with their BPS bound states, reproduce the W-boson spectrum in the dual theory [9, 10, 11, 12]. With the electric charge turned on, dyonic bound states built from these monopoles realize the $SL(2,\mathbb{Z})$ S-duality orbit of the dual W-bosons [4, 13].
@@
Our approach may be viewed as a smooth-monopole analogue of the Kapustin-Witten construction for geometric Langlands program [41], where ’t Hooft operators in the topologically twisted $\mathcal{N}=4$ theory are defined by imposing supersymmetric singular monopole behavior, and the resulting magnetic data are encoded in the geometry of Hecke modifications. In this framework, a ’t Hooft (Hecke) operator $T(w^{\vee})$ is labeled by a dominant weight $w^{\vee}$ of the Langlands dual group $G^{\vee}$, and the moduli space of singular monopoles inserted at a single point is identified with the space $\mathscr{Y}(w^{\vee})$ of Hecke modifications of type $w^{\vee}$ for a $G$-bundle. Its natural compactification $\overline{\mathscr{Y}}(w^{\vee})$ is a Schubert variety in the affine Grassmannian. The space of physical states $\mathcal{H}(w^{\vee})$ is given by the (intersection) cohomology of $\overline{\mathscr{Y}}(w^{\vee})$ and is isomorphic to $R(w^{\vee})$, the irreducible $G^{\vee}$-representation of highest weight $w^{\vee}$. Here we consider smooth monopoles arising from symmetry breaking, but the underlying geometric picture is similar. In particular, compactifying the Hecke modification space $\mathscr{Y}(w^{\vee})$ by adding all associated lower-weight Hecke modifications (which share the same topological type determined by $w^{\vee}$) mirrors the stratification of the monopole moduli space by allowing gauge-inequivalent magnetic charges at fixed topological charge.
@@
For example, when $SU(N+1)$ breaks to $U(N)$, there are NN $SU(2)$-embedded monopole solutions of unit topological charge, related by global $SU(N)$ Weyl transformations [33]. They are the classical counterparts of dual W-bosons in the fundamental of $SU(N)$. With (1.2) relaxed to (1.4), these solutions can be incorporated into a single connected moduli space labeled by the topological charge. The low-energy effective action for monopoles is obtained by a collective-coordinate expansion of $\mathcal{N}=4$ SYM. Starting from an arbitrary BPS trajectory, integrating out $A_{0}$ automatically enforces the background-gauge condition on the moduli derivatives. We will argue that, even with the background gauge imposed, the non-Abelian modes that do not commute with $G_{0}$ still have divergent norm. This divergence is also required for the stratified structure of the moduli space. For instance, in a relative moduli space consisting of an open stratum $\mathbb{C}^{2N}$ and a closed stratum $\mathbb{CP}^{2N-1}$, $\mathbb{CP}^{2N-1}$ must sit at the asymptotic boundary of $\mathbb{C}^{2N}$ and thus inherits a divergent metric. With the non-Abelian modes included, the moduli space is not necessarily hyper-Kähler. To be self-contained, we will derive the effective action via a collective-coordinate expansion without relying on hyper-Kähler structure [^2] The hyper-Kähler structure is needed only for the fermionic zero-mode counting, where it guarantees that the number of fermionic zero modes is twice the number of bosonic moduli. Here we take this relation as an input. In fact, with the non-normalizable modes included, the index calculation [21, 22] yields a moduli-space dimension that agrees with the dimension in the stratified formulation [38, 39]. Although this dimension is not necessarily divisible by $4$, the index still gives a $2:1$ ratio of fermionic to bosonic modes for $\mathcal{N}=4$ SYM [43, 44].. The resulting action takes the same form as in the maximally broken case, but the associated supersymmetry is $\mathcal{N}=2$ (enhanced to $\mathcal{N}=4$ when the moduli space is Kähler), rather than the $\mathcal{N}=8$ available for a hyper-Kähler target [42]. This reduction does not affect the physical conclusions: the $16$-fold degeneracy of the $\mathcal{N}=4$ massive vector multiplet is already saturated by the eight fermionic zero modes associated with the center-of-mass motion, while the ground states in the relative sector are still given by harmonic forms on the moduli space.
@@
Aside from the numerical match, the monopole ground states, which furnish the same $(H^{\vee})^{s}$-representation as the dual W-bosons, also carry definite weights of that representation (up to Weyl transformations). This allows us to construct magnetic operators generating the $(H^{\vee})^{s}$-action on the monopole Hilbert space. While the $H^{s}$-action is realized geometrically by isometric diffeomorphisms of the relative moduli space, the $(H^{\vee})^{s}$-action is realized algebraically on differential forms by wedge and contraction operations. In [25], it was noted that the harmonic forms on the coset space, which represent monopole ground states in that case, are $H^{s}$-invariant. Here we show that the ground states are $H^{s}$-invariant in general. Consequently, the $(H^{\vee})^{s}$-action may be implemented by operators commuting with both $H^{s}$ and the Hamiltonian. For monopole moduli spaces with a single stratum, we give an explicit construction of the resulting $H^{s}\times(H^{\vee})^{s}$-representation on the Hilbert space, thereby realizing the $H\times H^{\vee}$ GNO conjecture at the level of monopole quantum mechanics.
@@
The rest of the paper is organized as follows. Section 2 reviews the stratified structure of the monopole moduli space. Section 3 presents a general proof of the matching between the monopole (and dyon) spectrum and the dual W-boson spectrum when the unbroken gauge group is non-Abelian. In Section 4 we study the supersymmetric quantum mechanics on the monopole moduli space with the non-Abelian degrees of freedom included. In Section 5 we construct the magnetic gauge transformation generators and give an explicit realization of the $H^{s}\times(H^{\vee})^{s}$-representation on the monopole Hilbert space. We end with a discussion in Section 6.
@@ 2 Stratification of the monopole moduli space
Let $\bm{\alpha}_{a}^{\vee}=2\bm{\alpha}_{a}/(\bm{\alpha}_{a}\cdot\bm{\alpha}_{a})$ be the coroots. The unbroken gauge group is $H=H^{s}\times U(1)^{t}$, where $H^{s}$ is semisimple with simple roots $\bm{\alpha}_{t+1},\ldots,\bm{\alpha}_{r}$.
@@
$m_{a}$ are gauge invariant, while $q_{a}$ transform under the Weyl group of $H$. For each $1\leq i\leq n$, the intersection $\mathfrak{h}_{0}\cap C(k_{i})$ is a Weyl ($H$)-orbit. We can choose a unique representative by imposing the anti-dominant condition
@@
    $\mathcal{M}_{\mathrm{rel}}(1,\Phi_{0})=\mathcal{M}_{\mathrm{rel},1}(1,\Phi_{0})=C(k_{1})\cong\mathbb{CP}^{1}.$ (2.25)
@@
    The stratum $\mathcal{M}_{\mathrm{rel},2}(2,\Phi_{0})$ contains a parameter $a\in[0,\infty)$ describing the size of the non-Abelian cloud. The corresponding monopole solutions and the moduli space metric can be found in [45, 46, 47]. When $a\to\infty$, $G_{0}$ jumps from $k_{2}$ to $C(k_{1})$, consistent with the connectedness of $\mathcal{M}_{\mathrm{rel}}(2,\Phi_{0})$. Note that the boundary of $\mathcal{M}_{\mathrm{rel},2}(2,\Phi_{0})$ would be $7$-dimensional, while $\mathcal{M}_{\mathrm{rel},1}(2,\Phi_{0})$ is $6$-dimensional, so $\mathcal{M}_{\mathrm{rel}}(2,\Phi_{0})$ is not obtained by attaching a boundary to $\mathcal{M}_{\mathrm{rel},2}(2,\Phi_{0})$, but rather by compactifying $\mathcal{M}_{\mathrm{rel},2}(2,\Phi_{0})$ with $\mathcal{M}_{\mathrm{rel},1}(2,\Phi_{0})$ placed at infinity, analogous to compactifying $\mathbb{R}^{2}$ to $\mathbb{S}^{2}$ by adding a point at infinity.
@@ 3.2 $\mathcal{M}_{\mathrm{rel}}(1,\Phi_{0})=C(k_{1})\cup\mathcal{M}_{\mathrm{rel},2}(1,\Phi_{0})$
    where $S$ is the tautological rank-$2$ subbundle on $\mathrm{Gr}(2,N{+}1)$ and $\pi$ is the bundle projection. Gluing this open bundle to the closed Schubert variety at infinity yields a compactification [48]
@@ 3.3 General proof
For a compact, connected semisimple group $\mathcal{G}$, let $\lambda$ be an anti-dominant coweight of $\mathcal{G}$ (equivalently, a weight of $\mathcal{G}^{\vee}$), and let $V_{\lambda}$ be the irreducible $\mathcal{G}^{\vee}$-representation with lowest weight $\lambda$ [^5] Representations are usually labeled by a dominant highest weight; here we use the equivalent convention of an anti-dominant lowest weight to match our setting.. The set of weights in $V_{\lambda}$ decomposes into Weyl orbits
@@
      $H^{\ast}[\mathcal{M}_{\mathrm{rel},i}(1,\Phi_{0});\mathbb{R}]=H^{\ast}[C(k_{i});\mathbb{R}],$ (3.68)
@@
      $\sum_{k}\text{dim}H^{k}[\mathcal{M}_{\mathrm{rel},i}(1,\Phi_{0});\mathbb{R}]=\chi[\mathcal{M}_{\mathrm{rel},i}(1,\Phi_{0})]=\chi[C(k_{i})].$ (3.69)
@@
    In the full moduli space $\mathcal{M}(m,\Phi_{0})$, when $m>1$, each fiber $\mathcal{M}(m,\Phi_{0},k_{i})$ contains $m$ identical massive monopoles (with some additional massless monopoles when $i\neq 1$) and thus has the structure [4, 24]
@@
    In fact, when $m>1$ with $\mathcal{M}_{\mathrm{rel}}(m,\Phi_{0})$ noncompact, bound states are normalizable harmonic forms, which are typically not captured by the ordinary de Rham cohomology. In our situation, the metric scale along $C(k_{i})$ is infinite, so the usual normalizability criterion must be adapted to the fiber-base decomposition. Let $N_{\mathcal{M}}$ denote the number of bound states on $\mathcal{M}$. Instead of (3.75), a suitable counting is
@@
    and all topological charges become holomorphic except for mm. [^6] In the $SU(3)$ example 2 for $m=2$, take $\tilde{\Phi}_{0}=\mathop{\mathrm{diag}}\nolimits(2v,-v+\varepsilon,-v-\varepsilon)$ with $0<\varepsilon<3v$. Both $\mathcal{M}_{\mathrm{rel}}(\tilde{\Phi}_{0},k_{1})$ and $\mathcal{M}_{\mathrm{rel}}(2,\Phi_{0},k_{1})$ are the relative moduli space of two massive monopoles of identical charge $\mathop{\mathrm{diag}}\nolimits(1,-1,0)$, whereas $\mathcal{M}_{\mathrm{rel}}(\tilde{\Phi}_{0},k_{2})$ contains an additional massive monopole of charge $\mathop{\mathrm{diag}}\nolimits(0,1,-1)$. As $\varepsilon\to 0$, the $\mathop{\mathrm{diag}}\nolimits(0,1,-1)$ sector becomes massless and form the non-Abelian cloud in $\mathcal{M}_{\mathrm{rel}}(2,\Phi_{0},k_{2})$ [24]. Here no $\mathbf{k}_{i}$ is a root, so $N_{\mathcal{M}_{\mathrm{rel}}(2,\Phi_{0})}=0$. We may expect [^7] When $\tilde{\Phi}_{0}\rightarrow\Phi_{0}$, harmonic forms on $\mathcal{M}_{\mathrm{rel}}(\tilde{\Phi}_{0},k_{i})$ descend to $\mathcal{M}_{\mathrm{rel}}(m,\Phi_{0},k_{i})$. This limiting procedure has subtleties, see [24]. We will return to these issues in Example 3.3.
@@
    In the maximally broken regime, matching the monopole spectrum to the dual W-bosons requires